# The Dawn of Natural Language to SQL: Are We Fully Ready?

### v1

@@ 4.3. Experiments on Evaluating Efficiency
Finding 9. Based on the ratio of Execution Accuracy (EX) to the Average Cost per nl2sql task, we observe that prompt-based LLM methods calling GPT-3.5-turbo offer higher cost-effectiveness. Although DAILSQL(SC) shows EX improvements over DAILSQL on Spider and BIRD datasets, it introduces higher costs reducing its cost-effectiveness.
@@
Exp-7: Efficiency of PLM-based Methods. In practical applications, it is essential to consider both the performance and efficiency of nl2sql methods, including latency per sample. Different methods have varying hardware requirements, particularly GPU memory, which tends to increase with model size. Selecting the appropriate nl2sql method based on available hardware resources and latency requirements is a common challenge. We assess three metrics across six models: RESDSQL-Base/Large/3B and RESDSQL-Base/Large/3B + NatSQL, focusing on Execution Accuracy (EX), Latency Per Sample, and GPU Memory Used, utilizing the Spider development set for evaluation. Note that since model efficiency is dataset-agnostic, we omit the experiments on BIRD dataset due to space limitations.
Table 6 shows that as the model parameter size increases, so do the required GPU memory and latency for the same method. However, we find that RESDSQL-Base+NatSQL with 220M parameters and RESDSQL-Large with 770M parameters achieve similar EX scores (80.2% and 80.1%, respectively), with the former exhibiting lower latency per sample and requiring less GPU memory. A similar observation can be made when comparing RESDSQL-Large+NatSQL with RESDSQL-3B. Therefore, although different models may have similar EX scores, they can vary significantly in latency and hardware requirements. In practical scenarios, selecting an appropriate model should be based on latency requirements and available hardware resources.
@@
Table 7 reports experimental results. The highest VES score is highlighted in orange in the table. The methods with best VES on subsets of different difficulties are not consistent, and there is no clear winner between LLM-based and PLM-based methods. For the same method, it tends to have lower VES on more difficult subsets, possibly due to the increased complexity of sqls and the associated prediction challenge and execution time.
Finding 11. Based on VES metric, there is no clear winner between LLM-based and PLM-based methods. For the same method, it tends to have lower VES on more difficult subsets.
@@
Exp-9: The Impact of the #-Training Samples. In real-world scenarios, limited in-domain data often hinders performance. We conduct experiments on the Spider training set, randomly sampling subsets with size increments of 1000, as well as a smaller subset of size 500. Different methods are trained on these subsets, and their EX performance is evaluated on the Spider development set. The training hyper-parameters of RESDSQL-3B and RESDSQL-3B+NatSQL are the same with  (Li et al., 2023d), and the other methods are consistent with Exp-5.
@@ 5.1. A Design Space Exploration
(2) Prompting Strategy: Prompting strategies fall into zero-shot, where no nl2sql examples are included in the model input, and few-shot, which incorporates such examples, denoted as “3-shot”, “5-shot”, and etc., depending on the number of examples used. Table 1 shows PLM-based methods typically use zero-shot, while LLM-based methods vary: C3SQL (Dong et al., 2023) employs zero-shot, whereas DAILSQL(Gao et al., 2023) and DINSQL(Pourreza and Rafiei, 2023) use few-shot. The few-shot examples for DINSQL are manually designed and fixed, whereas those for DAILSQL are dynamically selected based on the similarity between the target question and training set examples.
@@
(c) Intermediate Representation strategy explores if a model employs an intermediary query form to address the natural language to sql translation’s mismatch problem, where sql’s design for relational databases doesn’t directly correlate with natural language semantics. Various solutions like  (Guo et al., 2019) and NatSQL (Gan et al., 2021) have been introduced. As shown in Table 1, models like LLM-based DINSQL (Pourreza and Rafiei, 2023) and several PLM-based methods (Li et al., 2023d; Rai et al., 2023; Gan et al., 2021) adopt NatSQL. In our setting, we only include NatSQL for simplification.
@@
(b) Self-Consistency involves executing various valid sql queries for a single nl query, using a voting mechanism on the outcomes to determine the most consistent sql as the final choice. It is used in C3SQL (Dong et al., 2023) and DAILSQL (Gao et al., 2023).
@@
(d) N-best Rerankers rank multiple candidate sql queries to select the most probable one as the final query (Zeng et al., 2023).
@@ 5.2. NL2SQL360 Facilitates Better NL2SQL
After categorizing different methods into a unified modular framework, it became clear that different methods use or propose new modules (still within our unified workflow) to enhance the performance of nl2sql solutions. This raises a question: Is it possible to achieve stronger performance by combining different modules from different nl2sql systems?
@@
Figure 14. NLSQL360-AAS Algorithm Overview.
@@
Step-4: nl2sql Module Mutation. Next, the individual $M^{s}_{t,i}$ (similarly $M^{s}_{t,j}$) will undergo module mutation in each layer (e.g., Pre-Processing Layer) based on the module mutation probability $p_{m}$. For example, if the Pre-Processing Layer of $M^{s}_{t,i}$ does not use the DB Contents module, a successful mutation will result in the inclusion of this module. After mutation, the individual is labeled as $M_{t+1,i}$ and will enter the next generation population. We repeat Steps 2, 3, and 4 until we obtain the complete next generation population $\{M_{t+1,n}\}^{N}_{n=1}$, marking one entire population iteration.
5.3. A Case Study of NL2SQL360-AAS
@@
In this section, we validate the effectiveness of the NL2SQL360-AAS algorithm. The search space is defined as shown in Figure 13. Note that, for simplification, we only utilize the Few-shot module from DAILSQL in the Prompting Strategy. Additionally, since we use GPT as our backbone, we cannot control the model’s decoding behavior, thus we only employ Greedy Search in the Decoding Strategy. We use the Spider development set as the target dataset and Execution Accuracy (EX) as the target nl2sql metric. The population size $N$ is set to 10, the number of population generations $T$ is 20, and the probabilities for nl2sql module swap and nl2sql module mutation, $p_{s}$ and $p_{m}$, are set to 0.5 and 0.2, respectively. To save on costs, we use GPT-3.5-turbo as the backbone model.
SuperSQL. As a result, from the final generation produced by the algorithm, we select the individual with the highest Execution Accuracy as our final searched nl2sql solution, namely SuperSQL. We find that the composition of SuperSQL is as follows: (1) in the Pre-Processing layer, it utilizes the Schema Linking module from RESDSQL and the DB Contents module from BRIDGE v2; (2) in the Prompting layer, it uses DAILSQL’s Few-shot module that dynamically selects in-context examples based on similarity; (3) in the SQL Generation layer, it uses OpenAI’s default Greedy-decoding strategy, excluding Multi-step or NatSQL intermediate representations; (4) in the Post-Processing layer, it incorporates the Self-Consistency module from DAILSQL. We explored the organization of the prompt for this composition, as illustrated in Figure 15. Under this combination, based on the DAILSQL prompt, the RESDSQL schema linking module is used to filter out irrelevant schema items. Furthermore, it incorporates the DB content module from the BRIDGE v2 method, employing string-matching algorithms to match the nl query with content in the database. Relevant content is then added as comments following the corresponding columns in the prompt, thereby enriching the information about the columns. Then, we replace the backbone model with GPT-4 for more powerful performance.
From the final generation produced by NL2SQL360-AAS, we select the individual with the highest EX metric as our final nl2sql solution, i.e., SuperSQL.
The Effectiveness of SuperSQL. We evaluate SuperSQL on the Spider development set, achieving 87.0% in EX and outperforming other competitive methods (Table 3). For sql queries of varying hardness, SuperSQL achieves the best results in Medium, Hard, and Extra hardness level subsets, demonstrating its effectiveness. Additionally, on the BIRD development set, SuperSQL shows competitive performance (Table 4).
We also evaluate SuperSQL on the Spider and BIRD test sets. SuperSQL achieves 87.0% EX on the Spider test set, ranking 2nd on the leaderboard, and 62.66% EX on the BIRD test set, ranking ninth. Note that SuperSQL surpasses all baselines within its design space. Specifically, SuperSQL outperforms the strongest baseline—DAILSQL(SC)—by 5.25% in the EX metric on the BIRD test set. This improvement is primarily due to our NL2SQL360-AAS, which effectively searches for superior module combinations based on different baselines in the design space. We expect that including more powerful baselines in the design space will further enhance our SuperSQL systems through NL2SQL360-AAS.
@@ 6. Research Opportunities
Interpret NL2SQL Solution. (i) NL2SQL Debugger can detect incorrect sql queries and allows users to step through the sql generation process, identify errors or mismatches, and understand the logic behind the generated sql. (ii) SQL and Query Results Interpretation method helps users understand whether the generated sql and query results meet their requirements.
@@
Adaptive Training Data Generation. The effectiveness of nl2sql methods depends greatly on the quality and coverage of training data. These methods often struggle with adapting to unseen databases. A promising research direction is the automatic generation of (nl, sql) pairs based on the model evaluation feedback. The key idea is that we dynamically synthesize (nl, sql) pairs, addressing both the challenge of domain adaptation and the need for high-quality, diverse training data, by utilizing insights gained from nl2sql performance evaluations.

### v2

@@ 4.3. Experiments on Evaluating Efficiency
Finding 9. Based on the ratio of Execution Accuracy (EX) to the Average Cost, we observe that prompt-based LLM methods calling GPT-3.5-turbo offer higher cost-effectiveness. Although DAILSQL(SC) shows EX improvements over DAILSQL on Spider and BIRD datasets, it introduces higher costs reducing its cost-effectiveness.
@@
Exp-7: Efficiency of PLM-based Methods. In practical applications, considering both the performance and efficiency of nl2sql methods is essential, especially latency per sample. Different methods have varying GPU memory requirements, increasing with model size. Selecting the appropriate method based on available hardware and latency requirements is a common challenge. We assess three metrics across six models: RESDSQL-Base/Large/3B and RESDSQL-Base/Large/3B + NatSQL, focusing on Execution Accuracy (EX), Latency Per Sample, and GPU Memory Used, using the Spider development set. Model efficiency is dataset-agnostic, so we omit the BIRD dataset due to space constraints.
Table 6 shows that as model size increases, so do GPU memory and latency. However, RESDSQL-Base+NatSQL (220M parameters) and RESDSQL-Large (770M parameters) achieve similar EX (80.2% and 80.1%), with the former having lower latency and memory usage. Similarly, RESDSQL-Large+NatSQL and RESDSQL-3B have comparable EX but differ in latency and hardware needs. Therefore, model selection should consider latency and hardware resources.
@@
Table 7 reports experimental results. The highest VES score is highlighted in orange in the table. The methods with the highest VES on subsets of varying difficulty are inconsistent, with no clear advantage for LLM-based or PLM-based approaches. Generally, a method shows lower VES on harder subsets, likely due to increased complexity and longer execution time.
Finding 11. Based on the VES metric, there is no clear winner between LLM-based and PLM-based methods. For the same method, it tends to have lower VES on more difficult subsets.
@@
Exp-9: The Impact of the #-Training Samples. In real-world scenarios, limited in-domain data often hinders performance. We experiment on the Spider training set, sampling subsets with size increments of 1000 and a smaller subset of 500. We train various methods on these subsets and evaluate their EX performance on the Spider development set. The training hyper-parameters for RESDSQL-3B and RESDSQL-3B+NatSQL match those in  (Li et al., 2023d), while other methods follow Exp-5.
@@ 5.1. A Design Space Exploration
(2) Prompting Strategy: Prompting strategies fall into zero-shot, where no nl2sql examples are included in the model input, and few-shot, which incorporates such examples, denoted as “3-shot”, “5-shot”, etc., depending on the number of examples used. Table 1 shows PLM-based methods typically use zero-shot, while LLM-based methods vary: C3SQL (Dong et al., 2023) employs zero-shot, whereas DAILSQL(Gao et al., 2024) and DINSQL(Pourreza and Rafiei, 2023) use few-shot. The few-shot examples for DINSQL are manually designed and fixed, whereas those for DAILSQL are dynamically selected based on the similarity between the target question and training set examples.
@@
(c) Intermediate Representation strategy explores if a model employs an intermediary query form to address the nl to sql translation’s mismatch problem, where sql’s design for relational databases doesn’t align with natural language semantics. Various solutions like  (Guo et al., 2019) and NatSQL (Gan et al., 2021) have been introduced. LLM-based DINSQL (Pourreza and Rafiei, 2023) and several PLM-based methods (Li et al., 2023d; Rai et al., 2023; Gan et al., 2021) adopt NatSQL. In our setting, we only include NatSQL for simplification.
@@
(b) Self-Consistency involves executing various valid sql queries for a single nl query, using a voting mechanism on the outcomes to determine the most consistent sql as the final choice. It is used in C3SQL (Dong et al., 2023) and DAILSQL (Gao et al., 2024).
@@
(d) N-best Rerankers rank multiple candidate sql queries to select the most probable one as the final query (Zeng et al., 2022).
@@ 5.2. NL2SQL360 Facilitates Better NL2SQL
After categorizing different LLM- and PLM-based methods into a unified modular framework, a question arises: Can automatically exploring and combining different modules from various nl2sql solutions achieve stronger performance?
@@
Figure 14. NL2SQL360-AAS Algorithm Overview.
@@
Step-4: nl2sql Module Mutation. Next, the individual $M^{s}_{t,i}$ (similarly $M^{s}_{t,j}$) will undergo module mutation in each layer (e.g., Pre-Processing Layer) based on the module mutation probability $p_{m}$. For example, if the Pre-Processing Layer of $M^{s}_{t,i}$ does not use the DB Contents module, a successful mutation will result in the inclusion of this module. After mutation, the individual is labeled as $M_{t+1,i}$ and will enter the next generation population.
We repeat Step 2–4 until we obtain the complete next generation population $\{M_{t+1,n}\}^{N}_{n=1}$, marking one entire population iteration.
5.3. SuperSQL: A Superior NL2SQL Solution Suggested by NL2SQL360-AAS
@@
In this section, we validate the effectiveness of the NL2SQL360-AAS algorithm. The search space is defined as shown in Figure 13. For simplification, we utilize only the Few-shot module from DAILSQL in the Prompting Strategy. Since we use GPT as our backbone, we cannot adjust its decoding strategy and thus we can only use Greedy Search in the Decoding phase. To save costs, we use GPT-3.5-turbo in the algorithm. The population size $N$ is set to 10, the number of population generations $T$ is 20, and the probabilities for nl2sql module swap and nl2sql module mutation, $p_{s}$ and $p_{m}$, are set to 0.5 and 0.2, respectively. We run the algorithm on the Spider development set with the EX as the target metric.
SuperSQL. We select the individual with the highest Execution Accuracy from the final generation produced by the NL2SQL360-AAS as our final searched nl2sql solution, namely SuperSQL.
The composition of SuperSQL is as follows: (1) Pre-Processing: Schema Linking from RESDSQL and DB Contents from BRIDGE v2; (2) Prompting: DAILSQL’s Few-shot module selects in-context examples by similarity; (3) SQL Generation: OpenAI’s Greedy-decoding strategy, without Multi-step or NatSQL; (4) Post-Processing: Self-Consistency from DAILSQL. We explored the organization of the prompt for this composition, as illustrated in Figure 15. Under this combination, the RESDSQL schema linking module is used to filter out irrelevant schema items. Furthermore, it incorporates the DB content module from the BRIDGE v2 method. The relevant content is added as comments next to the columns in the prompt, enhancing column information. We then replace the backbone model with GPT-4 for improved performance.
The Effectiveness of SuperSQL. We evaluate SuperSQL on the Spider development set, achieving 87.0% in EX and outperforming other methods (Table 3). SuperSQL excels in Medium, Hard, and Extra subsets, proving its effectiveness. On the BIRD development set, SuperSQL also shows competitive performance (Table 4).
We also evaluate SuperSQL on the Spider and BIRD test sets, achieving 87.0% EX on Spider (2nd place) and 62.66% EX on BIRD (9th place). Note that SuperSQL surpasses all baselines within its design space. Specifically, SuperSQL outperforms the strongest baseline—DAILSQL(SC)—by 5.25% on the BIRD test set. This improvement is primarily due to our NL2SQL360-AAS, which effectively searches for superior module combinations based on different baselines in the design space. Adding more powerful baselines is expected to further improve SuperSQL through NL2SQL360-AAS.
@@ 6. Research Opportunities
Interpret NL2SQL Solution. (i) NL2SQL Debugger detects incorrect sql queries, enables users to step through the sql generation process, and helps identify errors or mismatches. (ii) SQL and Query Results Interpretation method helps users assess if the generated sql and query results meet their requirements.
@@
Adaptive Training Data Generation. The effectiveness of nl2sql methods depends greatly on the quality and coverage of training data. These methods often struggle with adapting to unseen databases. A promising research direction is dynamically generating (nl, sql) pairs using model evaluation feedback. This approach tackles domain adaptation and ensures diverse, high-quality training data by leveraging insights from nl2sql performance.
@@ 7. Conclusion
Acknowledgements.
This paper was supported by Guangdong Basic and Applied Basic Research Foundation (2023A1515110545), National Key R&D Program of China (2023YFB4503600), NSF of China (61925205, 62232009, 62102215), Zhongguancun Lab, Huawei, TAL education, Beijing National Research Center for Information Science and Technology (BNRist), CCF-Huawei Populus Grove Fund (CCF-HuaweiDB202306), and the Fundamental Research Funds for the Central Universities.